# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v2

@@ IV-E2 PARC: Avoid Set with Tracking Error
Follows from Propositions 15 and Proposition 9. ∎
@@
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and from (15), the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ of the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\tilde{\mathcal{O}}_{t,i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$. Define the intermediate avoid set ${\Lambda}_{i,t,t}\subset{\Omega}_{t}$ as:
${\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{other}}},\tilde{\mathcal{O}}_{t,i}\right)\cap{\Omega}_{t}.$ (62)
@@ V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), and tight approximation of planning model reachable sets (Section V-C). We also assess the utility of learning-based reach-avoid in Sections V-A and V-B.
Key result: PARC outperforms other methods in goal reaching, safety, and computation time, but requires careful choice of planning model.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [15] and 3-D [5], and a more general quadrotor model in 3-D [10]. Implementation details are reported in Appendix -E. All experiments, and training of learning-based methods, are run on a desktop computer with a 24 core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
@@ V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deploy PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [15] approach.
@@
which are held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [31].
@@
V-A1 Experiment Setup
We use the obstacle and goal environment from [15, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 4 and Fig. 5).
We use PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piece-wise affinize (78) with 80 PWA regions, timestep $\Delta t=0.1$unitsec, and final time $t_{\mathrm{\textnormal{f}}}=2$unitsec. We assess the computation time, as well as the size of the BRAS to investigate the impact of planning model design.
@@ V-A4 Discussion
In Appendix -G and the next few examples, we show how a careful design of the planning model would lead to much better performance on more complicated scenarios, as well as enable stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
@@ V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [15] learning-based reach-avoid approach. For this experiment, we use a 3D near-hover quadrotor introduced for FaSTrack [5], a reachability-based planner tracker that uses adversarial planning and tracking models.
@@ V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -E. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.5 m) to reach its goal as shown in Fig. 7. We simulate 8,100 trials from different initial conditions. We measure the number of trials that reach the goal, and the number that crashes. We also consider the total computation time of each method.
@@ V-B2 Hypotheses
We anticipate that our PARC planner always reaches the goal safely. We also anticipate faster computation time of our reachability analysis versus the other two approaches, but note that FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides formal proof of safety, we anticipate it has no crashes. However, since it treats the planner and tracker as adversarial, we expect it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expect it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
@@ V-B3 Results
Fig. 7 illustrates results for all methods with an initial position of $[4,-1,3]^{\intercal}$. FaSTrack (orange) fails to find a feasible plan that reaches the goal, the trajectory rolled out by Neural CLBF (magenta) violates safety, while the trajectory rolled out by PARC (blue) finds a plan (black) that reaches the goal safely despite tracking error.
Fig. 6 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF finds 576 successful trajectories (7.1 %), while safety is violated for the rest. FaSTrack finds 0 successful trajectories but never violates safety. However, PARC finds the plan to reach the goal without collision for 2,592 initial states (32 %).
@@
Figure 6: Simulated trajectories of 10D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [5], Neural CLBF [15]. The green cuboid is the goal set and the pink cuboids are obstacles.
Figure 7: Realized trajectory of the 10D quadrotor in the environment for 10 seconds using three methods: ours (blue), FaSTrack [5] (orange), Neural CLBF [15] (magenta). The black dashed line denotes the plan computed using PARC. The goal is represented as a green cube while the obstacle is magenta cuboids. The boundary of environments exist although not illustrated in the figure.
Figure 8: Comparison of tracking error function between Fastrack and PARC for planning horizon $t_{\mathrm{\textnormal{f}}}=10$unitsec. FaSTrack computes a fixed, worst-case tracking error bound for all time because it assumes an adversarial planner, whereas PARC takes advantage of the smaller, time-varying tracking error that results from a planner and controller working together. The pale pink trajectories are samples used to compute the error as in Section IV-D.
@@ V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ cuboid. This often results in the inability to compute feasible plans in spaces tighter than the TEB. However, as seen in the Fig. 8, sampled tracking error under parameterized plan, the tracking error is much smaller than TEB, enabling PARC to find a plan in challenging scenarios.
@@ V-C Impact of Tight Approximation for Planning Reachability
Finally, we seek to understand the importance of computing a tighter approximation of reachable set for the planning model. To do this, we compare against RTD implemented on a general 3-D quadrotor [10] using zonotope reachability [23]. This allows us to use the same tracking error model but employs a different approach to planning model reachability analysis.
@@ V-C1 Experiment Setup
Our experimental setup is similar to that of Section V-B but with a narrower gap.
We use the same planning model as quadrotor RTD [10], which generates time-varying polynomial positions in 3-D [38] with $\Delta t=0.02$unitsec with $t_{\mathrm{\textnormal{f}}}=3$unitsec; this formulation is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [10].
@@
We declare success if the quadrotor reaches the goal region within 30 seconds without colliding with any obstacles.
@@ V-C2 Hypothesis
We anticipate that PARC will always safely find a way to reach the goal since we know the narrow gap is large enough to do so, and PARC’s primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipate that RTD will reach the goal much less often because it suffers an additional course of conservativeness from using an approximate method to compute the reachable set of its planning model.
@@ V-C4 Discussion
V-D Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using Backward Reach-Avoid Set (BRAS). By careful use of PWA systems, plus bounded tracking error, we are able to generate safe motion plans in cases where other methods are either too conservative or fail to compute entirely.
@@ V-D Overall Discussion
However, PARC still has room for improvement. Most obviously, the safe plan computation time is on the order of seconds for just a few obstacles; this is too slow for real-time operation. Furthermore, the planning model requires careful hand-crafting as shown in our experiments.

### v3

@@ IV-E2 PARC: Avoid Set with Tracking Error
This follows directly from Proposition 18 and Proposition 11. ∎
@@
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and from (40), the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ of the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\tilde{\mathcal{O}}_{t,i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$. Define the intermediate avoid set ${\Lambda}_{i,t,t}\subset{\Omega}_{t}$ as:
${\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{nonETI}}},\tilde{\mathcal{O}}_{t,i}\right)\cap{\Omega}_{t}.$ (44)
@@ V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), tight approximation of planning model reachable sets (Section V-C), and respect of time-varying tracking error in planning level (Section V-D). We also assess the utility of learning-based reach-avoid in Sections V-A and V-B.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [dawson2022safe] and 3-D [chen2021fastrack], and a more general quadrotor model in 3-D [kousik2019safe]. Implementation details are reported in Appendix -E. All experiments, and training of learning-based methods, are run on a desktop computer with a 24 core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
Key result: PARC outperforms other methods in goal reaching and safety in near-danger scenarios, but requires careful choice of planning model. While it can still guarantee reach-avoid in non-near-danger scenarios, its performance is worse than existing methods. Most experiments are summarized in Table II.
@@ V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deploy PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [dawson2022safe] approach.
Key result: Naïve design of PARC’s planning model can impact performance negatively. We confirm that a better choice of planning model solves the same task with higher performance in Appendix -G.
V-A1 Experiment Setup
@@
which are held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [kong2021ilqr].
@@
We use the obstacle and goal environment from [dawson2022safe, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 5 and Fig. 6).
We use PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piece-wise affinize (54) with 80 PWA regions, timestep $\Delta t=0.1$units, and final time $t_{\mathrm{\textnormal{f}}}=2$units. We assess the computation time, as well as the size of the BRAS to investigate the impact of planning model design.
@@ V-A4 Discussion
In Appendix -G and the next few examples, we show how a careful design of the planning model leads to much better performance on more complicated scenarios, as well as enabling stabilizing behavior such as braking without explicitly modeling velocity as a goal state.
@@ V-B Impact of Planner-Tracker Cooperation
We now study cooperative vs. adversarial planning and tracking models. We also further study the Neural CLBF [dawson2022safe] learning-based reach-avoid approach. For this experiment, we use a 3D near-hover quadrotor introduced for FaSTrack [chen2021fastrack], a reachability-based planner tracker that uses adversarial planning and tracking models.
Key result: PARC takes advantage of cooperative planner-tracker behavior to reduce conservativeness.
@@ V-B1 Experiment Setup
The tracking model and PARC’s planning model are listed in Appendix -E. The quadrotor must navigate a narrow gap (width 1.8 m, compared to drone width 0.54 m) to reach its goal as shown in Fig. 7. We simulate 8,100 trials from different initial conditions. We measure the number of trials that reach the goal, and the number that crashes. We also consider the total computation time of each method. Further experiment details are in Appendix -F.
[…]
@@ V-B2 Hypotheses
We anticipate that our PARC planner always reach the goal safely. We also anticipate faster computation time of our reachability analysis versus the other two approaches, but note that FaSTrack does not need to recompute the reachable set for every obstacle. Since FaSTrack provides formal proof of safety, we anticipate it has no crashes. However, since it treats the planner and tracker as adversarial, we expect it to struggle to reach the goal. On the other hand, since CLBFs specifically balance safety and liveness, we expect it to reach the goal more often than FaSTrack but at the expense of crashing due to the high chance of learning an incorrect representation of safety when forced to operate near obstacles.
@@ V-B3 Results
The overall results are reported in Table II. Fig. 7 illustrates results for all methods with an initial position of $[4,-1,3]^{\intercal}$. FaSTrack (orange) fails to find a feasible plan that reaches the goal, the trajectory rolled out by Neural CLBF (pink) violates safety, while the trajectory rolled out by PARC (blue) finds a plan that reaches the goal safely despite tracking error.
Fig. 8 illustrates the trajectories simulated on the grid of 8,100 initial positions using each method. Neural CLBF found 576 successful trajectories (7.1 %), while safety was violated for the rest. FaSTrack found 0 successful trajectories but never violated safety. However, PARC found the plan to reach the goal without collision for 2,592 initial states (32 %).
@@
Figure 8: Simulated trajectories of 10-D quadrotor for 8,100 initial states using three methods: PARC, FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe]. The green box is the goal set and the pink boxes are obstacles.
Figure 9: Comparison of tracking error function between FaSTrack and PARC for planning horizon $t_{\mathrm{\textnormal{f}}}=10$units. FaSTrack computes a fixed, worst-case tracking error bound for all time because it assumes an adversarial planner, whereas PARC takes advantage of the smaller, time-varying tracking error that results from a planner and controller working together. The pale pink trajectories are samples used to compute the error as in Section IV-D.
@@ V-B4 Discussion
The performance of FaSTrack is notably compromised by the large, conservatively computed TEB, specified as a $2.47\times 2.47\times 0.1$ box. This often results in the inability to compute feasible plans in spaces tighter than the TEB. However, as seen in Fig. 9, sampled tracking error under parameterized plan, the tracking error is much smaller than TEB, enabling PARC to find a plan in challenging scenarios.
@@
TABLE II: Performance Evaluation of PARC and Other Reach-Avoid Frameworks: This table compares success and safety rates across various systems and scenarios. The success rate is calculated as the percentage of simulations achieving their goals without collisions, while the safety rate measures the avoidance of obstacles. Computation times are categorized into offline and online phases, with the offline phase not involving known goals, obstacles, or initial states. We report the mean and the standard deviation for time.
[…]
10D Quadrotor PARC 0.32 1.00 $15.34\pm 1.44$ 115 V-B
FaSTrack 0.00 1.00 Timeout 3,600
Neural CLBF 0.07 0.07 18,000 0
PARC-TEF 0.07 1.00 $65.42\pm 9.73$ 2,880 V-C
13D Quadrotor RTD 0.00 1.00 Timeout 2,880
(Narrow Gap) PARC-Funnel 0.08 1.00 $12.20\pm 1.70$ 0 V-D
KDF 0.00 1.00 Timeout 0
13D Quadrotor PARC-Funnel 1.00 1.00 $3.07\pm 0.06$ 00 V-D
(Wide Gap) KDF 1.00 1.00 $0.44\pm 0.27$ 0
@@ V-C Impact of Tight Approximation for Planning Reachability
We seek to understand the importance of computing a tighter approximation of the reachable set for the planning model. To do this, we compare against RTD implemented on a general 3-D quadrotor [kousik2019safe] using zonotope reachability [althoff2015introduction]. This allows us to use the same tracking error model but employs a different approach to planning model reachability analysis.
@@ V-C1 Experiment Setup
We use the same planning model as quadrotor RTD [kousik2019safe], which generates time-varying polynomial positions in 3-D [mueller2015computationally] with $\Delta t=0.02$units with $t_{\mathrm{\textnormal{f}}}=3$units; this formulation is affine time-varying in the trajectory parameters with no non-ETI states, so reachability can be tightly approximated by PARC. Both PARC and RTD use the same tracking error that is precomputed in [kousik2019safe].
[…]
We declare safe if the quadrotor does not crash, and success if the quadrotor reaches the goal region without colliding with any obstacles.
@@ V-C2 Hypothesis
We anticipate that PARC will always safely find a way to reach the goal since we know the narrow gap is large enough to do so, and PARC’s primary source of conservativeness is its tracking error estimation (which is identical to that of RTD). We anticipate that RTD will reach the goal much less often because it suffers an additional course of conservativeness from using a polynomial planning model with zonotopic reachable sets.
@@ V-C4 Discussion
V-D Impact of Time-Varying Tracking Error
Figure 11: Comparison of PARC (a-b) and KDF (c-d) motion planning for a 3-D quadrotor across two scenarios: a narrow gap of $0.85m$ (left) and a wide gap of $3m$ (right), projected onto the $p_{x}$-$p_{y}$ plane. The initial planning state is $\mathbf{p}_{0}=[4.8,1,4]^{{}^{\top}}$ with a planning horizon of $t_{\mathrm{\textnormal{f}}}=15$units. Goals are in the green box, the dense pink boxes indicate the actual obstacles and the lighter pink boxes show the obstacles buffered by the drone’s volume. Pink trajectories within blue volumes represent computed motion plans and bounds of realized trajectories when tracking the plan (i.e. funnels), illustrating how each method handles tracking errors at the motion-planning level.
Finally, we explore the importance of accounting for the temporal correlation of tracking errors. In this experiment, we compare our approach with the KDF framework that combines sampling-based planning with funnel-based feedback control [verginis2022kdf]. This control mechanism prescribes that tracking error remain in funnels described by a performance function, $\rho(t)$, which decays exponentially over time [bechlioulis2014low]. Employing the same funnel-based control across both the KDF and PARC methodologies allows us to maintain a consistent tracking error model. This setup confirms that any observed differences in performance are solely attributed to the planner’s strategy of using tracking errors.
Key result: By considering time-varying as opposed to worst-case tracking error, PARC significantly reduces conservativeness of motion planning. Furthermore, PARC’s use of parameterized trajectories enabled the design of a funnel controller that maintained the prescribed tracking error bounds, whereas the same controller could not reliably track the smoothed output of a sampling-based motion planner.
Remark 21.
As shown in the other experiments, PARC can leverage a data-driven tracking error function that is potentially less conservative than the prescribed $\rho(t)$. In this experiment, we use $\rho(t)$ to ensure a fair comparison between PARC and KDF.
V-D1 Experiment Setup
We extend the setup from Section V-B to compare PARC and KDF using gap widths of 0.85 m, and 3 m, relative to the quadrotor’s 0.54 m of body width. We evaluate success and safety ratios, along with computation times, by simulating the quadrotor from zero velocity at 3,700 and 2,400 uniformly distributed initial positions across each scenario.
For the planner, PARC employs the polynomial planning model from Section V-C, but with $t_{\mathrm{\textnormal{f}}}=15\sec$. $\Delta t$ is $0.2\sec$ for the narrow gap and $0.5\sec$ for the wide gap. KDF employs RRT [lavalle1998rapidly] as its motion planner.
For the tracker, both PARC and KDF use a specialized funnel controller adapted for an under-actuated 3-D quadrotor proposed in [lapandic2022robust], as the conventional KDF funnel-based feedback controller is unsuitable for such systems [lapandic2023kinodynamic]. This controller conforms to the performance function $\rho(t)=\rho_{0}+(\rho_{0}-\rho_{\infty})e^{-\lambda t}$ where we choose $\rho_{0}=0.5,\rho_{\infty}=0.1,\lambda=0.5$. We note that this controller does not respect input limits. The 48 hyperparameters for the funnel controllers are reported in Appendix -F.
Despite extensive tuning efforts, a universal hyperparameter set for KDF could not be determined that respected the prescribed tracking error bounds; thus, we assume the ideal performance of the funnel controller, and base our statistics on reference trajectories ignoring tracking error. In other words, the comparison is biased towards benefitting KDF and disadvantaging PARC. We note that we designed controllers that function for a selected few initial states with KDF, the outcomes of which are demonstrated in Appendix -F and Fig. 11. In contrast, since PARC uses parameterized reference trajectories, we were able to succesfully identify a universal set of hyperparameters that enabled us to base the statistics on realized trajectories that included tracking error.
V-D2 Hypothesis
In the wide-gap scenario, both PARC and KDF are hypothesized to successfully execute reach-avoid plans. We anticipate KDF to demonstrate faster computation times, benefiting from its sampling-based planning approach. In the narrow-gap scenario, KDF is expected to struggle due to its reliance on maximum funnel values for free space construction in RRT, whereas PARC is predicted to succeed by leveraging the time-varying nature of tracking errors. We expect no safety violations for all experiments.
V-D3 Results
[…]
V-D4 Discussion
As anticipated, PARC demonstrates superior performance over KDF in narrow-gap scenarios due to its ability to utilize the time-varying nature of tracking errors effectively; thus taking advantage of a low tracking-error regime. This advantage is visually depicted in Fig. 11 a, where PARC waits for the funnel controller to reduce the tracking error before navigating through narrow gaps. In contrast, for wider gaps shown in Fig. 11 b, the quadrotor proceeds immediately. Conversely, KDF struggles in narrow spaces, as depicted in Fig. 11 c, because its planner computes paths based on a static maximum tracking error that exceeds the actual gap width (i.e., $\rho_{0}=0.5>0.31$), but performs well in wider spaces where this limitation does not impede planning as depicted in Fig. 11 d.
Notably, the difficulty in tuning the funnel controller to follow the smoothed reference trajectory proposed by RRT has led us to assume perfect funnel control in reporting KDF’s statistics. This aligns with observations in Section V-B that parameterized trajectories foster better cooperation with controllers compared to trajectories derived from smoothing splines over RRT waypoints, which can create adversarial tracking conditions.
PARC achieves reach-avoid plans in narrow-gap scenarios but faces lengthy computation times due to detailed sampling from BRAS and smaller $\Delta t$ steps. In contrast, KDF uses sampling-based motion planning (SBMP) effectively in scenarios with adequate gap widths, swiftly creating viable plans. This indicates a potential research direction: integrating SBMP within PARC could shorten computation times and facilitate the generation of expert trajectories that adeptly handle near-danger scenarios, surpassing conventional kinodynamic SBMP methods such as KDF.
V-E Overall Discussion
Through the above experiments, we have shown how our proposed PARC approach confers advantages in computing and using the Backward Reach-Avoid Set (BRAS). In near-danger scenarios, PARC was less conservative than FaSTrack and KDF since the former uses a time-variant tracking error function, whereas the latter two assumes constant, worst-case tracking errors. While RTD can similarly handle time-variant tracking error functions, PARC still outperforms it due to the lower numerical approximation error from representing the planning model as a PWA system instead of a polynomial and using H-polytopes instead of zonotopes. Finally, Neural CLBF cannot provide hard reach-avoid guarantees, resulting in many crashes with obstacles.
@@ V-E Overall Discussion
However, PARC still has room for improvement. Most obviously, the safe plan computation time is on the order of seconds for just a few obstacles; this is too slow for real-time operation. In contrast, RTD [kousik2019safe], although more conservative and failed to reach the goal, can compute a new plan about every 51 ms in Section V-C. Typically, we expect motion planners with real-time capabilities to compute new plans for at least every 0.5 s [kousik2020bridging]. To achieve this for PARC, parallelization needs to be implemented as a future work. Furthermore, the planning model requires careful hand-crafting as shown in our experiments.